OpenStudy (anonymous):

I need to find the 2 points for 2x+2y=16 and y=x-4

OpenStudy (anonymous):

pick any x value, plug it into both of your equatiosn and solve for y, repeat the process to get another point

OpenStudy (anonymous):

I kinda need like a step by step if thats possible. then ill be able to do the rest.

OpenStudy (anonymous):

try using 0 for x in y = x-4 y=0-4 y=